OpenStudy (anonymous):

What is the most precise name for quadrilateral ABCD with vertices A(-4,-1), B(-2,-5), C(4,-2) and D(2,2)? Parallelogram rhombus quadrilateral rectangle

OpenStudy (anonymous):

@Hartnn

OpenStudy (anonymous):

So, I used the midpoint formula. M = x1 + x2/2, y1 + y2/2 AB M = (-1 + -2/2, -1 + -5/2) = (-1.5,-3) BC M = (-2 + 4/2, -5 + -2/2) = (1,-3.5) CD M = (4+ 2/2, -5 + -2/2) = (3,0) DA M = (2 + -4/2, 2 + -1/2) = (-1,0.5)